OpenStudy (anonymous):

i will give medal and become fan Use formulas to find the lateral area and surface area of the given prism. Round your answer to the nearest whole number. (1 point) 795 m2; 809 m2 758 m2; 849 m2 795 m2; 822 m2 758 m2; 822 m2

OpenStudy (anonymous):

I can't see the prism. Can you upload an image please?

OpenStudy (anonymous):

Lateral Area = perimeter of triangle * L = \[(3+9+9.49) \times 37 = 795\] Surface Area = Perimeter of Triangle * L + 2*Area of Triangle = \[(3+9+9.49) \times 37 = 795\] + \[2 \times 0.5 \times 3 \times 9 = 27\] = \[795+27 = 822\] Lateral Area = 795 Surface Area = 822

OpenStudy (anonymous):

So Answer C then.

OpenStudy (anonymous):

Thank you, is that really all you have to do for the lateral? I missed that lesson and i havent been able to figure it out

OpenStudy (anonymous):

Yeah. That's it. Good luck. That function is for any prism as well. So if you ever encounter a prism, you can use that formula to calculate the lateral area.